The perimeter of a rectangle is 82 cm. If the length of the rectangle is 6 more than 4 times the width, what are the dimensions of the rectangle?
Answer by soumilaggarwal2001@gmail.com(3)

Let the width be x.
The the length = 4x+6
The perimeter of a rectangle = 2(l+b)
Here it is equal to 2((4x+6)+x)
=2(5x+6)
=10x+12
Now, 10x+12=82
Then, 10x=82-12=70
This implies that, x=70/10=7cm
The width = x = 7cm
The length = 4x+6 = 4*7+6 = 28+6 = 34cm